# The Simons Observatory: Science Goals and Forecasts for the Enhanced Large Aperture Telescope

### v2

@@ 1 Introduction and overall goals
Expected instrumental and map-depth properties for the fully completed, nine-year SO LAT survey (2025-2034). Two sensitivity targets are presented (baseline and goal), as in Ade et al. (2019).
@@ 2 Expanded capabilities
The mapping speed for the wide survey with the SO LAT (Parshley et al. 2018; Gudmundsson et al. 2021) will be almost doubled by fully populating the LAT receiver (LATR; Zhu et al. 2021) with six additional optics tubes (OTs), thus adding roughly 30,000 detectors.222The initial seven OTs in the LATR comprise one low-frequency, four mid-frequency, and two ultra-high-frequency OTs (Ade et al. 2019). The six new OTs will comprise four mid-frequency (MF) OTs containing dichroic detectors with bands centered at 93 and 145 GHz (Sierra et al. 2025), and two ultra-high-frequency (UHF) OTs containing dichroic detectors with bands centered at 225 and 280 GHz. After installation of these new detectors, the receiver will contain eight MF, four UHF, and one low-frequency (LF) OTs, containing a total of roughly 60,000 detectors and fully populating the receiver’s available focal plane. The anticipated sensitivities of these detectors, including effects due to atmospheric noise, are described in Ade et al. (2019). The angular resolution of the LAT at these frequencies (e.g., FWHM 1.4 arcmin at 145 GHz) is also described there. Table 1 summarizes the frequency channels, resolution, sensitivity, and detector counts for the fully populated LATR. We note that these values reflect current best estimates derived from ongoing technical development of the SO LAT hardware, as well as the currently planned sky coverage of the SO LAT survey, which is wider than that assumed in Ade et al. (2019). Further discussion of the sensitivities can be found in Appendix A.
@@
These new components are planned to be complete by 2028 (see timeline in Fig. 1), while the initial SO survey is ongoing using the half-populated LATR (Haridas et al. 2024). Observations with the fully populated receiver will then commence, with a survey planned through 2034. The cumulative survey duration, including both the initial phase and the full-sensitivity phase, is expected to be nine years. Note that the mapping speed of the LATR during the initial SO survey is roughly half that of the instrument during the latter part of the survey, after the upgrade. The forecasts presented in Sec. 3 are based on the cumulative sensitivity of the entire survey with “goal” noise levels, accounting for the sensitivity improvement at the nominal-to-fully-populated LATR transition (see Appendix A for further details).333The forecasts presented here include only statistical errors; systematic error contributions will be refined in future, dedicated studies. In addition, Planck data are incorporated in all cosmological forecasts unless explicitly stated otherwise — Planck is useful for measuring large-scale modes to which SO is less sensitive due to atmospheric noise, particularly in total intensity. As a reference point, the depth of the co-added wide-area SO 93 and 145 GHz maps is expected to reach 2.8 Karcmin at the conclusion of the survey in 2034, assuming the “goal” noise levels in Table 1.444Including all frequency maps, the depth is expected to reach 2.6 Karcmin; the LF and UHF channels are predominantly of use in foreground mitigation, rather than raw CMB sensitivity. Our forecasting approach follows that explained in detail in Ade et al. (2019): we generate simulated sky maps of the microwave sky at all SO and Planck frequencies, including all relevant sky components in temperature and polarization (with noise), and then process these through a harmonic-space internal linear combination pipeline to obtain post-component-separation noise power spectra, which are then used in all scientific forecasts. The exceptions to this full-survey-integrated forecasting approach are the time-domain science cases.555The cosmic birefringence forecast in Sec. 3.3 also does not include Planck data. These include the analysis of AGN light curves in Sec. 3.5, and millimeter-wave transient detection in Sec. 3.10, which both rely on the instantaneous sensitivity and transient detection pipeline of the SO survey with the fully populated LATR, rather than the cumulative depth of the full nine-year integrated data set.
@@ 3.1 Constraining the properties of primordial perturbations
By measuring primordial fluctuations in the CMB over twice the range of angular scales probed by the Planck satellite, the full SO survey (in combination with Planck) will significantly improve characterization of the scale dependence, Gaussianity, and adiabatic nature of the primordial density fluctuations that are the signature of the dynamics of the first moments of the universe (Meerburg et al. 2019; Slosar et al. 2019b; Hanany et al. 2019). SO will halve the current error bar on the scalar perturbation spectral index , testing the near-scale-invariant prediction of inflation over a wider range of scales than accessible to Planck (Slosar et al. 2019b). SO will further test early-universe models by constraining the Gaussianity of the perturbations to via kinematic Sunyaev-Zel’dovich (kSZ) tomography (Smith et al. 2018; Münchmeyer et al. 2019), improving current constraints by a factor of five (Meerburg et al. 2019), and also by constraining primordial isocurvature perturbations.666Direct estimates of from the primary CMB bispectrum will also provide robust bounds, albeit with error bars roughly 2-3 times larger than those expected from kSZ tomography. The primary bispectra will also tightly constrain other shapes of non-Gaussianity (equilateral and orthogonal), with roughly a factor of two improvement over the current bounds from Planck. Note that achieving via kSZ tomography requires overlapping galaxy survey data, as will be available on the SO footprint from Rubin, Euclid, and other surveys. Constraints on primordial tensor perturbations, which require data from the SATs, will be described in a separate publication, which will also describe delensing forecasts enabled by the enhanced LAT infrastructure (see Namikawa et al. (2022) and Hertig et al. (2024) for delensing forecasts for the nominal SO survey). Together, these measurements will provide the most detailed constraints to date on the primordial power spectrum and early-universe physics.
@@ 3.2 A refined image of the earliest snapshot of the universe
The reported tension between local and cosmological measurements of the rate of expansion of the universe (as characterized by the Hubble constant, ; Verde et al. 2019; Di Valentino et al. 2021a; Breuval et al. 2024; Freedman et al. 2024) may be resolved by new particle physics models that can increase the Hubble constant while preserving the fit to current CMB power spectrum data (see, e.g., Di Valentino et al. 2021b, for a review). Models that alter the pre-recombination dynamics are of particular theoretical interest (Knox & Millea 2020), and generically imprint signatures in the CMB temperature and polarization on small scales. Models that accelerate the recombination process itself are also of interest in this context (e.g., Chiang & Slosar 2018; Schöneberg et al. 2022).777We note that detailed knowledge of the recombination process is crucial for the interpretation of CMB data (Planck Collaboration et al. 2016a). In this context, the refined recombination codes CosmoRec (Chluba & Thomas 2011) and HyRec (Ali-Haïmoud & Hirata 2011) are designed to reach sufficient accuracy for modeling standard recombination in the analysis of SO data. The full SO survey data will discriminate among such models (e.g., Smith et al. 2020; Hill et al. 2022; Galli et al. 2022; Lynch et al. 2024; Kou & Lewis 2025), or potentially make a detection. The SO data will also enable sensitive searches for interacting dark matter particles, ultra-light axions, cosmic strings, and primordial magnetic fields, as well as precision tests of Big Bang Nucleosynthesis (Gluscevic et al. 2019; Grin et al. 2019; Grohs et al. 2019).
@@ 3.3 Improving constraints on cosmic birefringence
The SO LAT survey will observe a significant fraction of the Galactic plane, which will allow us to test this methodology and derive an independent, significantly tighter constraint (Diego-Palazuelos et al. 2022, 2023). Using two detector splits for each of the six LAT frequency channels over the full survey region, and in the multipole range , the full SO survey will yield , more than a factor of two improvement over current error bars, if the methodology and assumptions of Eskilt & Komatsu (2022) are applied to these data. Note that this forecast uses only SO LAT data, i.e., no Planck data are included, in contrast to the other cosmological forecasts in this paper.888Note that noise is included in the SO LAT noise model here, as in all other forecasts. Polarized thermal dust in the Milky Way shows a positive parity-violating correlation (Planck Collaboration et al. 2020c; Weiland et al. 2020), limiting current birefringence constraints due to the need to account for this poorly understood, parity-violating foreground; our forecast here assumes this “intrinsic” foreground contribution to the observed correlation to be negligible, but this assumption may not hold (Clark et al. 2021). Efforts are ongoing to understand the mechanisms by which dust could produce this parity-violating signal (e.g., Huffenberger et al. 2020; Clark et al. 2021; Cukierman et al. 2023; Diego-Palazuelos et al. 2023; Vacher et al. 2023; Halal et al. 2024; Hervías-Caimapo et al. 2024a). SO and complementary surveys (e.g., with the Fred Young Submillimeter Telescope; CCAT-Prime Collaboration et al. 2023) will enable characterization of the dust signal at high resolution and sensitivity, allowing improved modeling and more robust constraints on birefringence. Should a robust signal be detected with SO, it will be important to account for lensing-induced smoothing of the power spectrum in order to derive unbiased constraints on BSM physics (Naokawa & Namikawa 2023).
@@ 3.4 A new large-scale view of dark matter, baryons, and galaxy clusters
Neutrino oscillation experiments show the three neutrino species have a total mass of at least eV, or at least eV if the hierarchy of particle masses is inverted (Particle Data Group Collaboration 2014). In combination with DESI baryon acoustic oscillation data (Adame et al. 2025), the maps enabled by the enhanced infrastructure described in Sec. 2 will provide a measurement with eV (Ade et al. 2019). Combined with future improved constraints on the reionization optical depth, (e.g., from LiteBIRD, Lee et al. 2019), the SO + DESI measurement is expected to improve to eV.999We also anticipate competitive constraints from SO itself via the patchy kinematic Sunyaev-Zel’dovich effect, as explained later in this section; although these will not reach the precision expected from LiteBIRD, they will also improve the SO-derived neutrino mass constraint over that obtained with Planck constraints on . Anticipated improvements on from CLASS (Harrington et al. 2016) will also help in this regard. Evidence for non-zero neutrino mass from cosmological data, and the resulting constraints on the hierarchy, will be of significant consequence to particle physics, as detailed in the 2014 and 2023 P5 reports (HEPAP Subcommittee Collaboration 2014, 2023).
@@
Characterizing the energy efficiency of feedback in galaxies, sourced by supernovae and active galactic nuclei, is a major focus of galaxy evolution studies. Feedback is also the dominant physical uncertainty in optical gravitational lensing analyses that seek to constrain dark energy. Combined with measurements of large-scale structure from DESI, Rubin, and Euclid, the SO maps of integrated electron pressure (Compton-) and momentum from the thermal and kinematic Sunyaev-Zel’dovich (SZ) effects (Zeldovich & Sunyaev 1969; Sunyaev & Zeldovich 1972, 1980) will uniquely measure the baryon content in galaxies, groups, and clusters – a quantity not measurable with optical telescopes – and will constrain the feedback efficiency to the few percent level. Fig. 4 shows simulated stacked images of the ionized gas density in the circumgalactic medium around DESI and Rubin galaxies,101010For DESI, we assume 2.9 million luminous red galaxies in this stack, measured on an overlapping sky area of 10,000 ; for Rubin, we assume 1.5 billion galaxies measured on 16,000 . as measured using the kSZ effect with the full SO survey data. These measurements will improve over the precision of recent data from ACT and BOSS (Schaan et al. 2021, also shown in Fig. 4) by nearly two orders of magnitude (Battaglia et al. 2017) (see also, e.g., Hadzhiyska et al. (2024) for recent kSZ measurements using ACT and DESI and Coulton et al. (2024a); Bleem et al. (2022); McCarthy & Hill (2024); Tanimura et al. (2022); Chandran et al. (2023); Planck Collaboration et al. (2016b) for state-of-the-art Compton- maps from ACT, SPT, and Planck data).
@@
The SO maps will complement 21-centimeter experiments that probe the epoch of reionization (e.g., HERA Collaboration et al. 2023; Tingay et al. 2013; Mellema et al. 2013) by distinguishing among different models for how the universe was heated by the first ionizing sources. The “patchy” kSZ signal imprinted in the CMB during this epoch, due to scattering off the newly freed electrons, will enable the duration of the epoch of reionization to be measured with an uncertainty of (about 60 million years; Alvarez et al. 2019). This signal, interpreted using astrophysical reionization modeling, will also allow a determination of the optical depth , independent of large-scale CMB -mode measurements (Alvarez et al. 2021). We forecast that SO can reach using this new method, which relies on a joint analysis of the kSZ 2- and 4-point functions (Alvarez et al. 2021; Ferraro & Smith 2018).111111For the purposes of this forecast, foregrounds are treated as Gaussian; future work will refine this methodology (MacCrann et al. 2024; Raghunathan et al. 2024). This forecast is within a factor of two of the cosmic-variance-limited error bar from the large-scale -mode polarization targeted by satellite missions such as LiteBIRD (LiteBIRD Collaboration et al. 2023).
@@
Using the unique redshift independence of the thermal SZ (tSZ) effect, SO will extend tSZ cluster detection into the epoch in which the first massive, virialized structures formed at . Thermal SZ surveys from ACT, SPT, and Planck have led the field in constructing clean, complete, nearly mass-selected cluster samples, with the latest catalogs comprising several thousand galaxy clusters in total (Hilton et al. 2021; SPT, DES Collaboration 2024; Planck Collaboration 2016). We forecast that the full SO survey will detect clusters, with redshifts from overlapping DESI, Rubin, Euclid, and SPHEREx data, as well as dedicated optical and near-IR observations. This forecast uses the multifrequency matched-filter methodology described in Madhavacheril et al. (2017) (following Herranz et al. 2002; Melin et al. 2006) assuming the gas pressure profile from Arnaud et al. (2010)121212The exact pressure profile parameters used are: , , , , and , with a hydrostatic mass bias used to set the normalization of the pressure–mass relation. with a cut on SNR used to define the cluster sample. SO will thus provide the community with a homogeneous, well-defined catalog for follow-up studies out to high redshifts, with clusters in the unique discovery space (Mantz et al. 2019), enabling a broad range of multiwavelength science, including X-ray analyses with eROSITA (Merloni et al. 2024). Mass estimates will be enabled by galaxy weak lensing data from Rubin and Euclid for clusters at and by SO’s own CMB lensing data at higher redshifts. SO measurements of high- clusters and proto-clusters will also provide crucial total flux constraints for high-resolution follow-up observations with interferometric facilities (e.g., Kitayama et al. (2023); van Marrewijk et al. (2024)).

### v3

@@ 1 Introduction and overall goals
Expected instrumental and map-depth properties for the fully completed, nine-year SO LAT survey (2025-2034). Two sensitivity targets are presented (baseline and goal), as in Ade et al. (2019). The values in parentheses in the first column represent the approximate passband width of each frequency channel.
@@ 2 Expanded capabilities
The mapping speed for the wide survey with the SO LAT (Parshley et al. 2018; Gudmundsson et al. 2021) will be almost doubled by fully populating the LAT receiver (LATR; Zhu et al. 2021) with six additional optics tubes (OTs), thus adding roughly 30,000 detectors.222The initial seven OTs in the LATR comprise one low-frequency, four mid-frequency, and two ultra-high-frequency OTs (Ade et al. 2019). The six new OTs will comprise four mid-frequency (MF) OTs containing dichroic detectors with bands centered at 93 and 145 GHz (Sierra et al. 2025), and two ultra-high-frequency (UHF) OTs containing dichroic detectors with bands centered at 225 and 280 GHz. After installation of these new detectors, the receiver will contain eight MF, four UHF, and one low-frequency (LF) OTs, containing a total of roughly 60,000 detectors and fully populating the receiver’s available focal plane. The anticipated sensitivities of these detectors, including effects due to atmospheric noise, are described in Ade et al. (2019). The angular resolution of the LAT at these frequencies (e.g., FWHM 1.4 arcmin at 145 GHz) is also described there. Table 1 summarizes the frequency channels, resolution, sensitivity, and detector counts for the fully populated LATR. We note that these values reflect current best estimates derived from ongoing technical development of the SO LAT hardware, as well as the currently planned sky coverage of the SO LAT survey, which is wider than that assumed in Ade et al. (2019). Estimates of the bandwidth of each frequency channel are also provided, with each edge determined by the point at which the response drops to 50% of the peak. These are computed from the simulated optical coupling and on-chip filters. Further discussion of the sensitivities can be found in Appendix A, including the per-OT noise-equivalent temperatures (NETs).
@@
These new components are planned to be complete by 2028 (see timeline in Fig. 1), while the initial SO survey is ongoing using the half-populated LATR (Haridas et al. 2024). Observations with the fully populated receiver will then commence, with a survey planned through 2034. The cumulative survey duration, including both the initial phase and the full-sensitivity phase, is expected to be nine years. Note that the mapping speed of the LATR during the initial SO survey is roughly half that of the instrument during the latter part of the survey, after the upgrade. The forecasts presented in Sec. 3 are based on the cumulative sensitivity of the entire survey with “goal” noise levels, accounting for the sensitivity improvement at the nominal-to-fully-populated LATR transition (see Appendix A for further details). The forecasts presented here include only statistical errors; systematic error contributions will be refined in future, dedicated studies. In addition, Planck data are incorporated in all cosmological forecasts unless explicitly stated otherwise — Planck is useful for measuring large-scale modes to which SO is less sensitive due to atmospheric noise, particularly in total intensity. As a reference point, the depth of the co-added wide-area SO 93 and 145 GHz maps is expected to reach 2.8 Karcmin at the conclusion of the survey in 2034, assuming the “goal” noise levels in Table 1.333Including all frequency maps, the depth is expected to reach 2.6 Karcmin; the LF and UHF channels are predominantly of use in foreground mitigation, rather than raw CMB sensitivity. Our forecasting approach follows that explained in detail in Ade et al. (2019): we generate simulated sky maps of the microwave sky at all SO and Planck frequencies, including all relevant sky components in temperature and polarization (with noise), and then process these through a harmonic-space internal linear combination pipeline to obtain post-component-separation noise power spectra, which are then used in all scientific forecasts. The exceptions to this full-survey-integrated forecasting approach are the time-domain science cases.444The cosmic birefringence forecast in Sec. 3.3 also does not include Planck data. These include the analysis of AGN light curves in Sec. 3.5, and millimeter-wave transient detection in Sec. 3.10, which both rely on the instantaneous sensitivity and transient detection pipeline of the SO survey with the fully populated LATR, rather than the cumulative depth of the full nine-year integrated data set.
@@ 3.1 Constraining the properties of primordial perturbations
By measuring primordial fluctuations in the CMB over twice the range of angular scales probed by the Planck satellite, the full SO survey (in combination with Planck) will significantly improve characterization of the scale dependence, Gaussianity, and adiabatic nature of the primordial density fluctuations that are the signature of the dynamics of the first moments of the universe (Meerburg et al. 2019; Slosar et al. 2019b; Hanany et al. 2019). SO will halve the current error bar on the scalar perturbation spectral index , testing the near-scale-invariant prediction of inflation over a wider range of scales than accessible to Planck (Slosar et al. 2019b). SO will further test early-universe models by constraining the Gaussianity of the perturbations to via kinematic Sunyaev-Zel’dovich (kSZ) tomography (Smith et al. 2018; Münchmeyer et al. 2019), improving current constraints by a factor of five (Meerburg et al. 2019), and also by constraining primordial isocurvature perturbations.555Direct estimates of from the primary CMB bispectrum will also provide robust bounds, albeit with error bars roughly 2-3 times larger than those expected from kSZ tomography. The primary bispectra will also tightly constrain other shapes of non-Gaussianity (equilateral and orthogonal), with roughly a factor of two improvement over the current bounds from Planck. Note that achieving via kSZ tomography requires overlapping galaxy survey data, as will be available on the SO footprint from Rubin, Euclid, and other surveys. Constraints on primordial tensor perturbations, which require data from the SATs, will be described in a separate publication, which will also describe delensing forecasts enabled by the enhanced LAT infrastructure (see Namikawa et al. (2022) and Hertig et al. (2024) for delensing forecasts for the nominal SO survey). Together, these measurements will provide the most detailed constraints to date on the primordial power spectrum and early-universe physics.
@@ 3.2 A refined image of the earliest snapshot of the universe
The reported tension between local and cosmological measurements of the rate of expansion of the universe (as characterized by the Hubble constant, ; Verde et al. 2019; Di Valentino et al. 2021a; Breuval et al. 2024; Freedman et al. 2024) may be resolved by new particle physics models that can increase the Hubble constant while preserving the fit to current CMB power spectrum data (see, e.g., Di Valentino et al. 2021b, for a review). Models that alter the pre-recombination dynamics are of particular theoretical interest (Knox & Millea 2020), and generically imprint signatures in the CMB temperature and polarization on small scales. Models that accelerate the recombination process itself are also of interest in this context (e.g., Chiang & Slosar 2018; Schöneberg et al. 2022).666We note that detailed knowledge of the recombination process is crucial for the interpretation of CMB data (Planck Collaboration et al. 2016a). In this context, the refined recombination codes CosmoRec (Chluba & Thomas 2011) and HyRec (Ali-Ha¨ımoud & Hirata 2011) are designed to reach sufficient accuracy for modeling standard recombination in the analysis of SO data. The full SO survey data will discriminate among such models (e.g., Smith et al. 2020; Hill et al. 2022; Galli et al. 2022; Lynch et al. 2024; Kou & Lewis 2025), or potentially make a detection. The SO data will also enable sensitive searches for interacting dark matter particles, ultra-light axions, cosmic strings, and primordial magnetic fields, as well as precision tests of Big Bang Nucleosynthesis (Gluscevic et al. 2019; Grin et al. 2019; Grohs et al. 2019).
@@ 3.3 Improving constraints on cosmic birefringence
The SO LAT survey will observe a significant fraction of the Galactic plane, which will allow us to test this methodology and derive an independent, significantly tighter constraint (Diego-Palazuelos et al. 2022, 2023). Using two detector splits for each of the six LAT frequency channels over the full survey region, and in the multipole range , the full SO survey will yield , more than a factor of two improvement over current error bars, if the methodology and assumptions of Eskilt & Komatsu (2022) are applied to these data. Note that this forecast uses only SO LAT data, i.e., no Planck data are included, in contrast to the other cosmological forecasts in this paper.777Note that noise is included in the SO LAT noise model here, as in all other forecasts. Polarized thermal dust in the Milky Way shows a positive parity-violating correlation (Planck Collaboration et al. 2020c; Weiland et al. 2020), limiting current birefringence constraints due to the need to account for this poorly understood, parity-violating foreground; our forecast here assumes this “intrinsic” foreground contribution to the observed correlation to be negligible, but this assumption may not hold (Clark et al. 2021). Efforts are ongoing to understand the mechanisms by which dust could produce this parity-violating signal (e.g., Huffenberger et al. 2020; Clark et al. 2021; Cukierman et al. 2023; Diego-Palazuelos et al. 2023; Vacher et al. 2023; Halal et al. 2024; Herv´ıas-Caimapo et al. 2025). SO and complementary surveys (e.g., with the Fred Young Submillimeter Telescope; CCAT-Prime Collaboration et al. 2023) will enable characterization of the dust signal at high resolution and sensitivity, allowing improved modeling and more robust constraints on birefringence. Should a robust signal be detected with SO, it will be important to account for lensing-induced smoothing of the power spectrum in order to derive unbiased constraints on BSM physics (Naokawa & Namikawa 2023; Idicherian Lonappan 2025).
@@ 3.4 A new large-scale view of dark matter, baryons, and galaxy clusters
Neutrino oscillation experiments show the three neutrino species have a total mass of at least eV, or at least eV if the hierarchy of particle masses is inverted (Particle Data Group Collaboration 2014). In combination with DESI baryon acoustic oscillation data (Adame et al. 2025), the maps enabled by the enhanced infrastructure described in Sec. 2 will provide a measurement with eV (Ade et al. 2019). Combined with future improved constraints on the reionization optical depth, (e.g., from LiteBIRD, Lee et al. 2019), the SO + DESI measurement is expected to improve to eV.888We also anticipate competitive constraints from SO itself via the patchy kinematic Sunyaev-Zel’dovich effect, as explained later in this section; although these will not reach the precision expected from LiteBIRD, they will also improve the SO-derived neutrino mass constraint over that obtained with Planck constraints on . Anticipated improvements on from CLASS (Harrington et al. 2016) will also help in this regard. Evidence for non-zero neutrino mass from cosmological data, and the resulting constraints on the hierarchy, will be of significant consequence to particle physics, as detailed in the 2014 and 2023 P5 reports (HEPAP Subcommittee Collaboration 2014, 2023).
@@
Characterizing the energy efficiency of feedback in galaxies, sourced by supernovae and active galactic nuclei, is a major focus of galaxy evolution studies. Feedback is also the dominant physical uncertainty in optical gravitational lensing analyses that seek to constrain dark energy. Combined with measurements of large-scale structure from DESI, Rubin, and Euclid, the SO maps of integrated electron pressure (Compton-) and momentum from the thermal and kinematic Sunyaev-Zel’dovich (SZ) effects (Zeldovich & Sunyaev 1969; Sunyaev & Zeldovich 1972, 1980) will uniquely measure the baryon content in galaxies, groups, and clusters – a quantity not measurable with optical telescopes – and will constrain the feedback efficiency to the few percent level. Fig. 4 shows simulated stacked images of the ionized gas density in the circumgalactic medium around DESI and Rubin galaxies,999For DESI, we assume 2.9 million luminous red galaxies in this stack, measured on an overlapping sky area of 10,000 ; for Rubin, we assume 1.5 billion galaxies measured on 16,000 . as measured using the kSZ effect with the full SO survey data. These measurements will improve over the precision of recent data from ACT and BOSS (Schaan et al. 2021, also shown in Fig. 4) by nearly two orders of magnitude (Battaglia et al. 2017) (see also, e.g., Hadzhiyska et al. (2024) for recent kSZ measurements using ACT and DESI and Coulton et al. (2024a); Bleem et al. (2022); McCarthy & Hill (2024); Tanimura et al. (2022); Chandran et al. (2023); Planck Collaboration et al. (2016b) for state-of-the-art Compton- maps from ACT, SPT, and Planck data).
@@
The SO maps will complement 21-centimeter experiments that probe the epoch of reionization (e.g., HERA Collaboration et al. 2023; Tingay et al. 2013; Mellema et al. 2013) by distinguishing among different models for how the universe was heated by the first ionizing sources. The “patchy” kSZ signal imprinted in the CMB during this epoch, due to scattering off the newly freed electrons, will enable the duration of the epoch of reionization to be measured with an uncertainty of (about 60 million years; Alvarez et al. 2019). This signal, interpreted using astrophysical reionization modeling, will also allow a determination of the optical depth , independent of large-scale CMB -mode measurements (Alvarez et al. 2021). We forecast that SO can reach using this new method, which relies on a joint analysis of the kSZ 2- and 4-point functions (Alvarez et al. 2021; Ferraro & Smith 2018).101010For the purposes of this forecast, foregrounds are treated as Gaussian; future work will refine this methodology (MacCrann et al. 2024; Raghunathan et al. 2024). This forecast is within a factor of two of the cosmic-variance-limited error bar from the large-scale -mode polarization targeted by satellite missions such as LiteBIRD (LiteBIRD Collaboration et al. 2023).
@@
Using the unique redshift independence of the thermal SZ (tSZ) effect, SO will extend tSZ cluster detection into the epoch in which the first massive, virialized structures formed at . Thermal SZ surveys from ACT, SPT, and Planck have led the field in constructing clean, complete, nearly mass-selected cluster samples, with the latest catalogs comprising several thousand galaxy clusters in total (Hilton et al. 2021; SPT, DES Collaboration 2024; Planck Collaboration 2016). We forecast that the full SO survey will detect clusters, with redshifts from overlapping DESI, Rubin, Euclid, and SPHEREx data, as well as dedicated optical and near-IR observations. This forecast uses the multifrequency matched-filter methodology described in Madhavacheril et al. (2017) (following Herranz et al. 2002; Melin et al. 2006) assuming the gas pressure profile from Arnaud et al. (2010)111111The exact pressure profile parameters used are: , , , , and , with a hydrostatic mass bias used to set the normalization of the pressure–mass relation. with a cut on SNR used to define the cluster sample. SO will thus provide the community with a homogeneous, well-defined catalog for follow-up studies out to high redshifts, with clusters in the unique discovery space (Mantz et al. 2019), enabling a broad range of multiwavelength science, including X-ray analyses with eROSITA (Merloni et al. 2024). Mass estimates will be enabled by galaxy weak lensing data from Rubin and Euclid for clusters at and by SO’s own CMB lensing data at higher redshifts. SO measurements of high- clusters and proto-clusters will also provide crucial total flux constraints for high-resolution follow-up observations with interferometric facilities (e.g., Kitayama et al. (2023); van Marrewijk et al. (2024)).